OpenStudy (anonymous):

Rebekah purchased a sweater that was reduced by twenty-five percent (or one-fourth) of its original price. She also used a store coupon for five dollars off her already-reduced purchase. When she made the purchase, she paid the cashier nineteen dollars. What was the original price of the sweater?

OpenStudy (texaschic101):

s = cost of the sweater s - 0.25s - 5 = 19 0.75s - 5 = 19 0.75s = 19 + 5 0.75s = 24 s = 24/0.75 s = 32 check.. 32 - 0.25(32) - 5 = 19 32 - 8 - 5 = 19 32 - 13 = 19 19 = 19 (correct) original price of sweater (s) = $ 32